(a) How many bit strings of length 8 are there?
(b) How many bit strings of length 8 or less are there?
(Count the empty string of length zero also.)
(c) How many strings of 6 lower case English letters are there that have the letter x in them somewhere? Here strings may use the same letter more than once. (Hint: It might be easier to first count the strings that don't have an x in them.)
(d) A particular brand of shirt comes in 12 colors, has a male version and a female version, and comes in 2 sizes for each sex. How many different types of this shirt are made?
1
Expert's answer
2020-10-28T18:48:47-0400

(a) "2^8=256"

There are 256 bit strings of length 8.


(b)


"2^0+2^1+2^2+2^3+2^4+2^5+2^6+2^7+2^8="

"=2^9-1=511"

There are 511 bit strings of length 8 or less.


(c) There are 26 letters in the alphabet.

The number of 6 letter strings "=26^6"

The number of 6 letter strings without any x "=25^6"

The number of 6 letter strings with at least one x


"26^6-25^6=64775151"

There are 64,775,151 strings of 6 lower case English letters that have the letter x in them somewhere.


(d)

Color "=12"

Gender "=2"

Size"=2"

Use the Product Rule


"12\cdot2\cdot2=48"

48 different types of this shirt are made.